Given the following information of first year students of a particular school.
​​​​​​​
Age (Years)
23
24
26
28
29
Number of students
8




Probability
1.5a
2.5 a
0.4
0.25
0.1

Find
i. a
ii. Interquartile range
1
Expert's answer
2020-12-20T16:52:27-0500

i. In a probability distribution, the sum of all "p_i" must be equal to 1. So

"1.5 a +2.5a+0.4+0.25+0.1 = 1"

"4a +0.75=1"

"4a = 0.25"

"a=0.0625"

ii. Interquartile range, by definition, is "Q_3 - Q_1" or "x_{75} - x_{25}"

5 values are given: 23 24 26 28 29

Given an odd 2n+1 number of values

first quartile Q1 = median of the n smallest values

third quartile Q3 = median of the n largest values

In out case, n=2, so "Q_1 =23.5" (median of 23 and 24) and "Q_3 = 28.5" (median of 28 and 29).

"IQR = 28.5 - 23.5 = 5"